OpenStudy (kimberlyevens):

cos 5x ��������� cos 7x = 0 @perl

OpenStudy (perl):

Θ

OpenStudy (midhun.madhu1987):

cos A - cos B = -2 Sin (A+B)/2. Sin (A-B)/2

OpenStudy (perl):

cos(5x) - cos(7x) = 0 ?

OpenStudy (kimberlyevens):

yes @perl sorry

OpenStudy (midhun.madhu1987):

you can use this formula to simplify: cos A - cos B = -2 Sin (A+B)/2. Sin (A-B)/2 where A = 5x and B= 7x

OpenStudy (perl):

cos(5x) = cos(7x)

OpenStudy (perl):

take inverse cosine of both sides

OpenStudy (perl):

midhun, how does that substitution make it easier to solve for x ? (exactly)

OpenStudy (kimberlyevens):

i dont really know thats why i posted my question

OpenStudy (perl):

what are the directions, solve for x ? in degrees or radians

OpenStudy (midhun.madhu1987):

Taking inverse I am not sure about it.. so I chose this method...

OpenStudy (midhun.madhu1987):

@perl can u please check this link? http://openstudy.com/study#/updates/5466d220e4b05b0e1f1daae8

OpenStudy (perl):

oh i see

OpenStudy (perl):

then you used zero product property

OpenStudy (midhun.madhu1987):

But I dont know whether this can be done in a simpler way taking inverse as you said

OpenStudy (midhun.madhu1987):

yess

OpenStudy (perl):

right, we can take invercosine of both sides

OpenStudy (midhun.madhu1987):

can u just do that for me

OpenStudy (perl):

sure

OpenStudy (midhun.madhu1987):

btw @kimberlyevens... you understood that method?

OpenStudy (perl):

cos(5x) = cos(7x) 5x = 7x + 2pi*n 5x = -7x + 2pi*n

OpenStudy (midhun.madhu1987):

 what would be the next step?